# Attributing Response to Context: A Jensen–Shannon Divergence Driven Mechanistic Study of Context Attribution in Retrieval-Augmented Generation

### v2

@@ 4 Attributing Top Relevant Context Sentences via JSD
In this section, we introduce our ARC-JSD to identify the most relevant context sentences contributing to generated responses. We subsequently present empirical evaluations that demonstrate the effectiveness of ARC-JSD compared to the method proposed by Cohen-Wang et al.cohen2024contextcite , across multiple datasets and varying scales of RAG-LLMs.
@@ 4.1 Identifying Relevant Context via JSD
Following the assumption proposed by Cohen-Wang et al.cohen2024contextcite , the removal of context segments critical to generating a specific response $\mathcal{R}$ significantly impacts the probability distribution of that response. Conversely, the removal of less relevant context segments is expected to minimally affect the probability distribution of $\mathcal{R}$.
Unlike the approach by Cohen-Wang et al.cohen2024contextcite , which requires extensive sampling of ablated contexts for each $(\mathcal{C},\mathcal{Q})$ pair and training a surrogate model to learn context-response relationships, our proposed ARC-JSD method relies purely on inference in the Figure 1. Specifically, we compute the JSD between the response probability distributions conditioned on the full context $\mathcal{C}$ and on each context-ablated variant $\mathcal{C}_{\text{ABLATE}}(c_{i})$:
@@
where we use $\text{JSD}(c_{i})$ to aggregate the JSD score of each generated tokens $r_{j}$ from $\mathcal{R}$ when the context sentence $c_{i}$ is ablated from the context $\mathcal{C}$. By calculating JSD scores for all sentences in the context, we identify the most relevant context sentence $c_{i}$ by selecting the sentence based on the assumption about the significant impact of removing critical context segments:
$c_{i}=\arg\max_{c_{i}\in\mathcal{C}}\left(\{\text{JSD}(c_{i})\}_{i=1}^{|\mathcal{C}|}\right)$ (2)
@@ 4.2 Evaluation of Context Attribution Accuracy
To assess the efficacy of our ARC-JSD method, we conduct experiments on three widely recognised question-answering datasets commonly used in RAG studies: TyDi QA (clark2020tydi, ): a multilingual QA dataset using the entire Wikipedia articles as the external context (we only use the English part), Hotpot QA (yang2018hotpotqa, ): a multi-hop QA dataset requiring reasoning for questions based on multiple documents, and MuSiQue (trivedi2022musique, ): a high-quality multi-hop QA benchmark over Wikipedia that highlights minimal context and multiple valid reasoning paths to evaluate complex reasoning capabilities. Table 1 summarises the statistics of these datasets, where MuSiQue has the longest context input compared to others With the average length of context in sentences $|\mathcal{C}|=93.6$. Our evaluations involve four instruction-tuned LLMs of varying scales, namely Qwen2-1.5B-IT, Qwen2-7B-IT (yang2024qwen2, ), Gemma2-2B-IT, and Gemma2-9B-IT (team2024gemma, ). For each dataset, we randomly select up to 1,000 samples from their development sets. All models are evaluated in inference mode without further fine-tuning. The performance of our ARC-JSD method is benchmarked against the Contextcite proposed by Cohen-Wang et al.cohen2024contextcite (Appendix D includes more details).
Table 2: Context attribution accuracy of Contextcite baseline and our ARC-JSD, where the Contextcite needs $n+1=256$ ablation calls to achieve a low root mean squared error based on their work cohen2024contextcite . Compared to cohen2024contextcite , our ARC-JSD only needs $|\mathcal{C}|$ ablation calls, which significantly smaller than $n$ based on Table 1.
Models Datasets Qwen2 1.5B IT Qwen2 7B IT Gemma2 2B IT Gemma2 9B IT Time Complexity
Contextcite ($n=$256 calls) TyDi QA 77.5 75.1 70.7 76.4 $o(n+1)$
Hotpot QA 54.9 68.4 54.8 70.4
MuSiQue 54.5 60.3 51.8 60.9
ARC-JSD TyDi QA 84.1 80.0 76.6 81.8 $o(|\mathcal{C}|+1)$
Hotpot QA 71.1 82.3 67.0 79.4
MuSiQue 60.6 76.8 65.3 78.2
Table 2 presents the comparative results, clearly demonstrating that ARC-JSD consistently outperforms Contextcite across all datasets and LLM scales, yielding an average accuracy improvement of approximately 10.7%. Crucially, while Contextcite typically necessitates constructing $n+1=256$ randomly ablated context samples per response-context pair for surrogate model training to achieve a low root mean squared error, ARC-JSD requires only $|\mathcal{C}|$ ablations per pair, a number considerably smaller based on dataset statistics provided in Table 1. Consequently, our method offers substantial computational efficiency improvements, achieving up to 3-fold speedups (See Appendix F for details). In addition, we utilise GPT-4.1 mini as a judge to compare whether the generated responses of all RAG models are semantically equivalent to the corresponding gold answers from the datasets when context attribution is correct. The average accuracy is up to 99.3% (See Appendix E for details.)
@@ 5 Mechanistically Study RAG LLMs for Context Attribution
Figure 2: Following our proposed ARC-JSD framework, we apply JSD-based metric to internal components of RAG LLMs: (a) For each attention head or MLP output at each layer, we can calculate the probability distribution of the same response $\mathcal{R}$ conditioned on the same query $\mathcal{Q}$ with full context $\mathcal{C}$ and ablated context $\mathcal{C}_{\text{ABLATE}}(c_{\text{top-1}})$ by removing the top relevant context sentence based on § 4.1; (b) We can further locate top-$N$ relevant attention heads or MLPs which contribute the context attribution by ranking the collected JSD scores with a descending order.
@@ 5.1 Locating Relevant Attention Heads and MLPs
To better understand the internal mechanisms by which RAG LLMs attribute generated responses to their relevant context sentences, we systematically investigate the specific attention heads and multilayer perceptron (MLP) layers involved. Our method combines the ARC-JSD metric described previously (§ 4.1) with the Logit Lens (Nostalgebraist2020, ) to precisely quantify contributions from these internal model components.
@@
$\begin{split}J_{\text{Top-}N}(\text{Attn})&=\text{sort}\left(\{\text{JSD}_{\text{Attn}}^{\ell,h}\}_{\ell=0,h=0}^{L,H},\text{descending}\right)\\ J_{\text{Top-}N}(\text{MLP})&=\text{sort}\left(\{\text{JSD}_{\text{MLP}}^{\ell}\}_{\ell=0}^{L},\text{descending}\right)\end{split}$ (5)
@@ 5.2 Mechanistic Insights from Located Attention Heads and MLPs
Applying the methodology described in § 5.1, we conducted experiments across three benchmark datasets (see § 4.2) using various LLM scales. Figure 3 presents the distribution and JSD scores of attention heads identified as most relevant for context attribution in Qwen2-1.5B-Instruct on TyDi QA dataset. Our analysis reveals that the top attention heads contributing to context attribution predominantly reside in the higher layers. This observation holds across most datasets, partially corroborating earlier findings by Wu et al.wu2025retrieval , which indicated that retrieval-related attention heads are typically found in the intermediate and higher layers.
Figure 3: The average JSD score of attention heads and MLP of Qwen2-1.5B-IT on TyDi QA dataset across all layers. The deeper colour indicates larger JSD scores.
Notably, our work expands upon the NIAH setting explored by Wu et alwu2025retrieval by mechanistically evaluating attention heads and MLPs relevance through paraphrasing and contextual integration of RAG LLMs. This setting better reflects real-world RAG applications, where models rarely copy text exactly but instead synthesise and rephrase information from retrieved sources. Additional visualisations and distributions for another Qwen2-7B-IT and Gemma2 models across all datasets are provided in Appendix H. Similarly, Figure 3 illustrates that the intermediate and higher MLP layers also significantly contribute to context attribution. This pattern remains consistent across different datasets and model scales within the same LLM family. Corresponding detailed findings for Qwen2-7B-IT and Gemma2 models across all datasets are available in Appendix H.
@@ 6.1 Semantic Gains of Attention and MLPs for Context Attribution
Based on the introduction of internal mechanism of LLMs in the § 3 and full context $\mathcal{C}$ with query $\mathcal{Q}$ as model’s input, we further split the residual stream flow of each layer into three parts for each generated token $t_{i}$, i.e., pre-residual stream $\mathbf{x}_{i}^{\ell,\text{pre}}$, middle-residual stream $\mathbf{x}_{i}^{\ell,\text{mid}}$ and post-residual stream $\mathbf{x}_{i}^{\ell,\text{post}}$:
$\mathbf{x}_{i}^{\ell,\text{pre}}=\mathbf{x}_{i}^{\ell-1,\text{post}}\quad\mathbf{x}_{i}^{\ell,\text{mid}}=\mathbf{x}_{i}^{\ell,\text{pre}}+\mathbf{a}_{i}^{\ell}\quad\mathbf{x}_{i}^{\ell,\text{post}}=\mathbf{x}_{i}^{\ell,\text{mid}}+\mathbf{m}_{i}^{\ell}=\mathbf{x}_{i}^{\ell+1,\text{pre}}$ (6)
After applying the logit lens to $\mathbf{x}_{i}^{\ell,\text{pre}},\mathbf{x}_{i}^{\ell,\text{mid}}$ and $\mathbf{x}_{i}^{\ell,\text{post}}$ via the softmax, we will have the probability distribution of the generated token $t_{i}^{\ell,\text{pre}}$, $t_{i}^{\ell,\text{mid}}$ and $t_{i}^{\ell,\text{post}}$ for each layer, and then we will use greedy decoding to select the top-1 token with the highest probability:
$t_{i}^{\ell,\text{pre/mid/post}}=\arg\max_{t_{i}^{\ell,\text{pre/mid/post}}\in\mathcal{\mathcal{V}}}\left(\text{softmax}\left(\text{logit}(\mathbf{x}_{i}^{\ell,\text{pre/mid/post}})\right)\right)$ (7)
Consequently, we can project the selected token $t_{i}^{\ell,\text{pre/mid/post}}$ into the vocabulary embedding space via the unembedding matrix $W_{U}\in\mathbb{R}^{d\times|\mathcal{V}|}$:
$\mathbf{e}_{i}^{\ell,\text{pre/mid/post}}=W_{U}[:t_{i}^{\ell,\text{pre/mid/post}}]$ (8)
We can calculate the corresponding semantic gains $\Delta_{i}^{\ell,\text{Attn}}$ and $\Delta_{i}^{\ell,\text{MLP}}$ via attention and MLP modules using the cosine similarity difference with the generated response token embedding $\mathbf{e}_{i}=W_{U}[:r_{i}]$:
$\Delta_{i}^{\ell,\text{Attn}}=\cos(\mathbf{e}_{i}^{\ell,\text{mid}},\mathbf{e}_{i})-\cos(\mathbf{e}_{i}^{\ell,\text{pre}},\mathbf{e}_{i}),\quad\Delta_{i}^{\ell,\text{MLP}}=\cos(\mathbf{e}_{i}^{\ell,\text{post}},\mathbf{e}_{i})-\cos(\mathbf{e}_{i}^{\ell,\text{mid}},\mathbf{e}_{i})$ (9)
Finally, we will average across the entire generated responses $\mathcal{R}$ and calculate the semantic gains $\Delta^{\ell,\text{Attn}}$ and $\Delta^{\ell,\text{MLP}}$ for attention MLP of each layer, and collect and sort the semantic gains of attention and MLP from all layer with descending order:

### v4

@@ 4.1 Identifying Relevant Context via JSD
Following assumption proposed by cohen2024contextcite , removal of context segments critical to generating a specific response $\mathcal{R}$ significantly impacts probability distribution of that response. Conversely, the removal of less relevant context segments is expected to minimally affect probability distribution of $\mathcal{R}$.
Unlike the approach by cohen2024contextcite , which requires extensive sampling of ablated contexts for each $(\mathcal{C},\mathcal{Q})$ pair and training a surrogate model to learn context-response relationships, our proposed ARC-JSD method relies purely on inference in the Fig. 1. Specifically, we compute the JSD between the response probability distributions conditioned on the full context $\mathcal{C}$ and on each context-ablated variant $\mathcal{C}_{\text{ABLATE}}(c_{i})$:
@@
where we use $\text{JSD}(c_{i})$ to aggregate the JSD score of each generated tokens $r_{j}$ from $\mathcal{R}$ when the context sentence $c_{i}$ is ablated from the context $\mathcal{C}$. By calculating JSD scores for all sentences in the context, we identify the most relevant context sentence $c_{i}$ by selecting the sentence based on the assumption about the significant impact of removing critical context segments: $c_{\text{Top-1}}=\arg\max_{c_{i}\in\mathcal{C}}\left(\{\text{JSD}(c_{i})\}_{i=1}^{|\mathcal{C}|}\right)$.
@@ 4.2 Evaluation of Context Attribution Accuracy
Table 1: The FLOPs for each baseline and ARC-JSD, where $P$ indicates the number of target model parameters, $T$ indicates the number of tokens per context sentence, $L$ is layer numbers of target model, $|\mathcal{R}|$ and $|\mathcal{C}|$ indicates the number of response tokens and context sentences, respectively.
Baselines Theoretical FLOPs Slowdown over ARC-JSD
ALTI-Logit $2PT|\mathcal{C}||\mathcal{R}|L$ $|\mathcal{R}|L/|\mathcal{C}|$
MIRAGE $4PT|\mathcal{C}|(2|\mathcal{C}|+1)$ $4+2/|\mathcal{C}|$
Contextcite (32 calls) $2PT\times 32^{2}$ $(32/|\mathcal{C}|)^{2}$
Contextcite (256 calls) $2PT\times 256^{2}$ $(256/|\mathcal{C}|)^{2}$
ARC-JSD (key-value caching) $2PT|\mathcal{C}|^{2}$ 1
To assess efficacy of our ARC-JSD method, we conduct experiments on three widely recognised question-answering datasets commonly used in RAG studies: TyDi QA (clark2020tydi, ): a multilingual QA dataset using entire Wikipedia articles as external context (we only use English part), Hotpot QA (yang2018hotpotqa, ): a multi-hop QA dataset requiring reasoning for questions based on multiple documents, and MuSiQue (trivedi2022musique, ): a high-quality multi-hop QA benchmark over Wikipedia that highlights minimal context and multiple valid reasoning paths to evaluate complex reasoning capabilities. Moreover, we evaluate our ARC-JSD with different training-free baselines for context attribution: ALTI-Logit (ferrando2023explaining, ): a method to directly compare logit difference between input context and generation on token level by accumulating layerwise logit; MIRAGE (qi2024model, ): a gradient-based and token-level method to locate context-sensitive tokens using contrastive feature attribution; Contextcite (cohen2024contextcite, ): a post-hoc method to train a linear surrogate model based on a fixed group of context ablation forward runs. Table 2 summarises the statistics of these datasets, where MuSiQue has the longest context input compared to others with the average length of context in sentences $|\mathcal{C}|=93.6$. Our evaluations involve four instruction-tuned LLMs of varying scales, namely Qwen2-1.5B-IT, Qwen2-7B-IT (yang2024qwen2, ), Gemma2-2B-IT, and Gemma2-9B-IT (team2024gemma, ). For each dataset, we randomly select up to 1,000 samples from their development sets. All models are evaluated in inference mode without further fine-tuning. We mainly evaluate the top-1 context attribution accuracy, which indicates the percentage of overlap between the predicted top-1 context sentence and gold-standard sentence on the datasets [^3] We choose sentence level because current QA datasets only have sentence-level gold labels to evaluate attribution accuracy. However, our ARC-JSD method can be extended to finer-grained interactions such as phrases or sub-sentences spans by dynamically selecting the start and end token indices.. For ALTI-Logit and MIRAGE, which mainly focus on token-level attribution, we use the accumulated operations to locate sentence-level context attribution prediction (Appendix E includes more details).
@@
Table 1 lists theoretical floating-point operations (FLOPs) for each method, where we follow the assumption from kaplan2020scaling ; hoffmann2022training , i.e., one forward pass needs approximately $2PT$ FLOPs. ARC-JSD is considerably cheaper than baselines because it pinpoints salient context sentences without back-propagation or iterative token masking. ContextCite requires a fixed 32 or 256 forward passes; this is economical only when input contains more than 32 or far more than 256 sentences, respectively, but its context attribution accuracy remains below that of ARC-JSD (see Fig. 2(a)). Fig. 2(a) presents compute-accuracy trade-off on MuSiQue dataset across all baselines and LLM backbones. It clearly demonstrates that ARC-JSD consistently outperforms all baselines, yielding an average context attribution accuracy improvement of approximately 10.7%. Although Contextcite-32 is more efficient when $|\mathcal{C}|$ is larger than 32, its attribution accuracy lags behind ARC-JSD. Overall, our method offers substantial computational efficiency improvements, achieving up to 3-fold speedups and consistently align with Pareto-optimal over multiple orders of magnitude for different LLM backbones. In addition, we utilise GPT-4.1 mini as a judge to compare whether generated responses of all RAG models are semantically equivalent to the corresponding gold answers from datasets when context attribution is correct. The average accuracy is up to 99.3% (See Appendix G and F for details).
@@ 5.1 Locating Relevant Attention Heads and MLPs
To better understand internal mechanisms by which RAG LLMs attribute generated responses to their relevant context sentences, we systematically investigate specific attention heads and MLP layers involved. Our method combines the ARC-JSD metric described previously (§ 4.1) with the Logit Lens (Nostalgebraist2020, ) to precisely quantify contributions from these internal model components.
@@
$J_{\text{Top-}N}(\text{Attn})=\text{sort}\left(\{\text{JSD}_{\text{Attn}}^{\ell,h}\}_{\ell=0,h=0}^{L,H},\text{descending}\right),J_{\text{Top-}N}(\text{MLP})=\text{sort}\left(\{\text{JSD}_{\text{MLP}}^{\ell}\}_{\ell=0}^{L},\text{descending}\right)$ (4)
Figure 2: (a) The compute-accuracy trade-off on MuSiQue for 4 baselines and ARC-JSD on 4 LLM backbones with GFLOPs $\mathrm{log}_{10}$ scale per sample; (b) The average JSD score of attention heads and MLP of Qwen2-1.5B-IT on TyDi QA across all layers. The deeper colour indicates larger JSD scores.
@@ 5.2 Mechanistic Insights from Located Attention Heads and MLPs
Applying the methodology described in § 5.1, we conducted experiments across three benchmark datasets (see § 4.2) using various LLM scales. Fig. 2(b) presents the distribution and JSD scores of attention heads identified as most relevant for context attribution in Qwen2-1.5B-Instruct on TyDi QA dataset. Our analysis reveals that the top attention heads contributing to context attribution predominantly reside in the higher layers. This observation holds across most datasets, partially corroborating earlier findings by wu2025retrieval , which indicated that retrieval-related attention heads are typically found in the intermediate and higher layers. Notably, our work expands upon NIAH setting explored by wu2025retrieval by mechanistically evaluating attention heads and MLPs relevance through paraphrasing and contextual integration of RAG LLMs. This setting better reflects real-world RAG applications, where models rarely copy text exactly but instead synthesise and rephrase information from retrieved sources. Additional visualisations and distributions for another Qwen2-7B-IT and Gemma2 models across all datasets are provided in Appendix I. Similarly, Fig. 2(b) illustrates that intermediate and higher MLP layers also significantly contribute to context attribution. This pattern remains consistent across different datasets and model scales within the same LLM family. Corresponding detailed findings for Qwen2-7B-IT and Gemma2 models are available in Appendix I.
@@ 6 Verification of JSD-based Mechanistic Study
Having identified critical RAG components using JSD in § 5.1, we now seek to verify these findings from a different analytical angle. Our JSD approach measures a component’s dependence on context through ablation. As a crucial cross-check, we introduce the semantic gain metric, which measures a component’s direct contribution to the correct answer when the full context is present. This allows us to test a key hypothesis: components that are truly important for attribution should not only depend on the context (high JSD) but also actively use it to improve the answer (high semantic gain). Demonstrating a strong correlation between these metrics will robustly verify that the components we located are indeed the primary drivers of context attribution.
@@ 6.1 Semantic Gains of Attention and MLPs for Context Attribution
Based on the introduction of internal mechanism of LLMs in the § 3 and full context $\mathcal{C}$ with query $\mathcal{Q}$ as model’s input, we further split the residual stream flow of each layer into three parts for each generated token $t_{i}$, i.e., pre-residual stream $\mathbf{x}_{i}^{\ell,\text{pre}}$, middle-residual stream $\mathbf{x}_{i}^{\ell,\text{mid}}$ and post-residual stream $\mathbf{x}_{i}^{\ell,\text{post}}$: $\mathbf{x}_{i}^{\ell,\text{pre}}=\mathbf{x}_{i}^{\ell-1,\text{post}},\mathbf{x}_{i}^{\ell,\text{mid}}=\mathbf{x}_{i}^{\ell,\text{pre}}+\mathbf{a}_{i}^{\ell},\mathbf{x}_{i}^{\ell,\text{post}}=\mathbf{x}_{i}^{\ell,\text{mid}}+\mathbf{m}_{i}^{\ell}=\mathbf{x}_{i}^{\ell+1,\text{pre}}$. After applying the logit lens to $\mathbf{x}_{i}^{\ell,\text{pre}},\mathbf{x}_{i}^{\ell,\text{mid}}$ and $\mathbf{x}_{i}^{\ell,\text{post}}$ via the softmax, we will have the probability distribution of the generated token $t_{i}^{\ell,\text{pre}}$, $t_{i}^{\ell,\text{mid}}$ and $t_{i}^{\ell,\text{post}}$ for each layer, and then we will use greedy decoding to select the top-1 token with the highest probability: $t_{i}^{\ell,\text{pre/mid/post}}=\arg\max_{t_{i}^{\ell,\text{pre/mid/post}}\in\mathcal{\mathcal{V}}}\left(\text{softmax}\left(\text{logit}(\mathbf{x}_{i}^{\ell,\text{pre/mid/post}})\right)\right)$. Consequently, we can project the selected token $t_{i}^{\ell,\text{pre/mid/post}}$ into the vocabulary embedding space via the unembedding matrix $W_{U}\in\mathbb{R}^{d\times|\mathcal{V}|}$: $\mathbf{e}_{i}^{\ell,\text{pre/mid/post}}=W_{U}[:t_{i}^{\ell,\text{pre/mid/post}}]$. We can calculate the corresponding semantic gains $\Delta_{i}^{\ell,\text{Attn}}$ and $\Delta_{i}^{\ell,\text{MLP}}$ via attention and MLP modules using the cosine similarity difference with the generated response token embedding $\mathbf{e}_{i}=W_{U}[:r_{i}]$: $\Delta_{i}^{\ell,\text{Attn}}=\cos(\mathbf{e}_{i}^{\ell,\text{mid}},\mathbf{e}_{i})-\cos(\mathbf{e}_{i}^{\ell,\text{pre}},\mathbf{e}_{i}),\Delta_{i}^{\ell,\text{MLP}}=\cos(\mathbf{e}_{i}^{\ell,\text{post}},\mathbf{e}_{i})-\cos(\mathbf{e}_{i}^{\ell,\text{mid}},\mathbf{e}_{i})$. Finally, we will average across the entire generated responses $\mathcal{R}$ and calculate the semantic gains $\Delta^{\ell,\text{Attn}}$ and $\Delta^{\ell,\text{MLP}}$ for attention MLP of each layer, and collect and sort the semantic gains of attention and MLP from all layer with descending order:
@@
Table 3: Spearman’s $\rho$ of the overlap about top-10 located attentions and MLPs between JSD-based mechanistic and semantic gain-based metrics over all datasets and RAG models. $\diamondsuit$ and $\spadesuit$ indicate $p$-value is << 0.05 and << 0.01, respectively.
Modules Top-1010 Datasets Qwen2 1.5B IT Qwen2 7B IT Gemma2 2B IT Gemma2 9B IT
[…]
@@ 6.2 Mutually Verifying JSD-based Mechanistic Study via the Semantic Gains of Attention and MLPs
where $\text{ranking of }(\cdot)$ will assign $1$ to the largest $\text{JSD}_{\text{MLP}}^{\ell}$ or $\Delta^{\ell,\text{MLP}}$ and the smallest $\text{JSD}_{\text{MLP}}^{\ell}$ or $\Delta^{\ell,\text{MLP}}$ will be assigned $L$. Then we uniform and remove the layer influence divided by $L$ to get $\text{ranking}_{J}$ and $\text{ranking}_{G}$, whose range is $[1/n,1]$, i.e., a smaller fraction will have a higher ranking ($1/n$ is best). Finally, we take the average of the $\text{ranking}_{J}$ and $\text{ranking}_{G}$ as the consensus $S^{(+)}$, where a smaller consensus inside of $S^{(+)}$ will indicate a stronger joint evidence that both metrics consider the layer important, and a larger consensus means at least one metric puts the layer far down the list. Finally, we calculate Spearman $\rho$ of $J_{\text{Top}-N}(\text{MLP})\cap S_{\text{Top}-N}^{(+)}$ and $G_{\text{Top}-N}(\text{MLP})\cap S_{\text{Top}-N}^{(+)}$, where $S_{\text{Top}-N}^{(+)}=\text{sort}(S^{(+)},\text{ascending})$. For attention components, we first average JSD scores of all attention heads in the same layer to build $\{\text{JSD}_{\text{Attn}}^{\ell}\}_{l=0}^{L}=\{\frac{1}{H}\sum_{h=0}^{H}\text{JSD}_{\text{Attn}}^{\ell,h}\}_{l=0}^{L}$, and then further calculate $\rho$ of $J_{\text{Top}-N}(\text{Attn})\cap S_{\text{Top}-N}^{(+)}$ and $G_{\text{Top}-N}(\text{Attn})\cap S_{\text{Top}-N}^{(+)}$. The benefit of using consensus $S^{(+)}$ instead of the raw JSD or semantic gain values is that $S^{(+)}$ will remove all scaling issue due to the different units and variances of JSD or semantic gains, and a single extremely large JSD or semantic gain will not swamp the fusion, which is robust to outliers.
Table 3 reports significant Spearman $\rho$ values for overlap between top-10 attention/MLP layers ranked by JSD and by semantic gain. This frequent co-occurrence indicates that both metrics track the same retrieval-driven signal that improves next-token prediction. Intuitively, when a layer genuinely draws on a retrieved sentence $c_{i}$ to write answer, ablating $c_{i}$ (i) alters that layer’s token distribution yielding high JSD and (ii) removes the “helpful push” toward the correct token lowering semantic gain. Layers that merely supply generic syntax or parametric knowledge may boost semantic gain without changing under ablation, so their JSD remains low; the strong overall correlation shows that such cases do not dominate, which further verifies effectiveness of ARC-JSD. In addition, ARC-JSD is practical: it requires only forward passes, avoiding cost and saturation issues of gradient-based saliency (qi2024model, ). Unlike KL (undefined with zero-probability bins) or logit-space $\ell_{2}$ distances (scale-dependent) (ferrando2023explaining, ), JSD is finite, symmetric, scale-free, and measured in interpretable bits.
7 Case Studies of Located Attention Heads and MLPs
Figure 3: The projection of $\mathbf{x}_{i}^{\ell,\text{mid}}$ and $\mathbf{x}_{i}^{\ell,\text{post}}$ via Logit Lens to vocabulary space from layer 20 to layer 27 of Qwen2-1.5B IT in TyDi QA data sample, where the generated response $\mathcal{R}$ is “A mosquito has two wings.” (See Appendix K for all layer projections). Each cell shows the most probable token decoded via Logit Lens. The colour indicates the probability of the decoded token of the corresponding $\mathbf{x}_{i}^{\ell,\text{mid}}$ or $\mathbf{x}_{i}^{\ell,\text{post}}$.
Visualisation of MLPs.
Based on semantic gains analysis from § 6.2, we visualise projection of middle-residual stream $\mathbf{x}_{i}^{\ell,\text{mid}}$ and post-residual stream $\mathbf{x}_{i}^{\ell,\text{post}}$ via Logit Lens to vocabulary space in Fig. 3 and Appendix K. In Fig. 3, Qwen2-1.5B-IT was given data from TyDi QA dev dataset with context about mosquitos introduction from Wikipedia and query “How many wings does a mosquito have?” as input, and it generates responses “A mosquito has two wings.” as output. Based on our ARC-JSD method, we successfully located top-relevant context sentence, i.e., “Mosquitoes have a slender segmented body, a pair of wings, three pairs of long hair-like legs, feathery antennae, and elongated mouthparts”. When we compare heatmap between $\mathbf{x}_{i}^{\ell,\text{post}}$ and $\mathbf{x}_{i}^{\ell,\text{mid}}$ in Fig. 3 from Layer 20 to Layer 27 (See Appendix K for full heatmap), we find that probability of correct token is increased significantly after $\mathbf{x}_{i}^{\ell,\text{post}}$ compared to $\mathbf{x}_{i}^{\ell,\text{mid}}$, such as ‘wings’ in Layer 23, ‘A’, ‘has’, ‘two’ in Layer 26, and ‘mosquito’, ‘two’, ‘A’ in Layer 27, which aligns with our findings that MLP contribute more parametric knowledge for context attribution in higher layers using JSD-based metric from § 5.2. Moreover, we find that several correct tokens are gradually transferred from their Chinese format to English version in Qwen2, such as ‘一只 (A)’, ‘拥有 (has)’ and ‘翅膀 (wings)’, which is reasonable as Chinese is one of main language resources used in Qwen2 pre- and post-training (yang2024qwen2, ). This finding also matches observations (wu2025the, ) that representations tend to be anchored by semantically-equivalent dominant-language tokens in higher layers.
Table 4: Hallucination rate and factual information maintenance by gating top-5 JSD-based attention heads and MLP using Qwen2 7B IT on Hotpot QA dataset. $\dagger$: Hallucination = RAG-generated response contradicts ground-truth while attributed context contained that ground-truth. $\ddagger$: we use JSD scores as a confidence gate to top-relevant attention heads and MLP layers with $\text{Mask}=0.7+0.3\times\text{sigmoid}(G)$, which means that top-relevant JSD attention and MLP will be multiplied with a mask close to 0.7 when their corresponding semantic gain $G$ is smaller than 0.
Setup for Qwen2 7B IT Hallucination Rate Pass@1 Factual F1 Notes
Base RAG 13.4% 76.1 One retrieved context sentence contains the answer
Gate Top-5 Attn & MLP 8.2% 75.9 Mask is 0.7 on top-5 JSD Attn & MLP when their $G<0$
Gate Random 5 Attn & MLP 12.7% 69.4 Control
Control RAGs based on ARC-JSD located Attention & MLPs.
Moreover, we conduct an ablation study to compare the JSD difference of responses by masking the top-10 relevant attention heads and randomly-selected 10 attention heads in Table 5. Generally, ablating attention heads located by using JSD-based metric causes larger JSD scores compared to the random attention heads ablation, which further verifies our proposed ARC-JSD can identify context-attribution-related attention heads.
Top-10 Attention Heads Random 10 Attention Heads